# PDLP: A Practical First-Order Method for Large-Scale Linear Programming

### v1

@@ Acknowledgements
We thank Ed Klotz for his helpful discussion regarding the Irish Electricity instance and Joey Huchette for his feedback on an early draft of the paper.
@@ Appendix A Reduced cost selection for PDLP
$r_{i}\leftarrow\begin{cases}c_{i}-a_{i}^{\top}y&\text{if }c_{i}-a_{i}^{\top}y>0\text{ and }x_{i}-(\ell_{v})_{i}\leq\absolutevalue{x_{i}}\\ c_{i}-a_{i}^{\top}y&\text{if }c_{i}-a_{i}^{\top}y<0\text{ and }(u_{v})_{i}-x_{i}\leq\absolutevalue{x_{i}}\\ 0&\text{otherwise}\end{cases}$
where $a_{i}$ is the $i$th column of $A$. This choice is more robust to artifically large bounds. For example, a relatively small residual reduced cost when a variable bound is extremely large, e.g., $10^{12}$ could prevent convergence through its effect on the dual objective value. This more advanced choice is only used when we run PDLP without feasibility polishing.
Appendix B Theoretical basis for feasibility polishing
For simplicity, we assume the LP instance takes the form $\min c^{\top}x$ such that $Ax=b,x\geq 0$ and focus on primal feasibility polishing (i.e., $c=\mathbf{0}$). However, this analysis generalizes to other settings such as the LP format described in Section 2 and dual feasibility polishing.
In this simple case, PDHG reduces to
@@ Appendix B Theoretical basis for feasibility polishing
Let $X_{F}:=\{x\geq 0:Ax=b\}$ be set of feasible solutions. Let $\mathop{\bf dist{}}(X,x):=\inf_{z\in X}\|x-z\|_{2}$. The Hoffman constant of $A$ is the largest constant, $H(A)$, such that
@@
Thanks to [48], we know that this constant is always positive.
@@
Let $Z^{\star}$ denote the set of minimax solutions to $\min_{x\geq 0}\max_{y}y^{\top}Ax-b^{\top}y$. From Corollary 2 of [8] we have
@@
Moreover, by [8] equation (22), we have
@@
where $+$ is the positive part operator. Combining (10), (13), (14), and (15) with this inequality yields (12). ∎
1
2Input: A starting point $x^{0,0}$, a step-size $\eta\in(0,\infty)$, a length for the restart period $\ell\in{\mathbb{N}}$;
3 Initialize outer loop counter $n\leftarrow 0$;
4 Initialize duals $y^{0,0}\leftarrow\mathbf{0}$;
5
6for $n=0,1,2,\dots$ do
7       initialize the inner loop. inner loop counter $t\leftarrow 0$ ;
8       repeat
9             Take a step of PDHG, i.e., calculate $(x^{n,t+1},y^{n,t+1})$ from $(x^{n,t},y^{n,t})$ via (9) ;
10             $\bar{x}^{n,t+1}\leftarrow\frac{1}{t+1}\sum_{i=1}^{t+1}x^{n,i}$;
11             $t\leftarrow t+1$;
12
13      until $t\geq\ell$ ;
14      restart the outer loop. $x^{n+1,0}\leftarrow\bar{x}^{n,t+1}$, $y^{n+1,0}\leftarrow\mathbf{0}$, $\tau^{n}\leftarrow\ell$, $n\leftarrow n+1$;
15
16 end for
@@ 11             $t\leftarrow t+1$;
$\|x^{k,0}-x^{0,0}\|_{2}$ $\leq\|x^{k,0}-x^{k-1,0}\|_{2}+\|x^{k-1,0}-x^{0,0}\|_{2}\leq(q+2)\mathop{\bf dist{}}(x^{k,0},X_{F})+\|x^{k-1,0}-x^{0,0}\|_{2}$ (17)
$\leq(q+2)2^{-k}\mathop{\bf dist{}}(x^{0,0},X_{F})+\|x^{k-1,0}-x^{0,0}\|_{2}$
@@
$\|x^{n,0}-x^{0,0}\|_{2}\leq\left(\sum_{i=1}^{n}2^{-i}\right)(q+2)\mathop{\bf dist{}}(x^{0,0},X_{F}).$
@@
In particular, (17) with $k=1$ shows the hypothesis for $n=1$ since $x_{k-1,0}=x^{0,0}$. If the hypothesis holds for $n=k-1$ then by (17) it holds for $n=k$. Thus we have proved the induction hypothesis. Using the formula for the sum of an infinite geometric series we get $\sum_{i=1}^{n}2^{-i}\leq\sum_{i=1}^{\infty}2^{-i}=1$, thus the induction hypothesis implies (16b). ∎

### v2

@@ Acknowledgements
We thank Ed Klotz for his helpful discussion regarding the Irish Electricity instance and his help clarifying the termination criteria of Gurobi. We also thank Joey Huchette for his feedback on an early draft of the paper.
@@ Appendix A Reduced cost selection for PDLP
[…]
where $a_{i}$ is the $i$th column of $A$. This choice is more robust to artificially large bounds. For example, a relatively small residual reduced cost when a variable bound is extremely large, e.g., $10^{12},$ could prevent convergence through its effect on the dual objective value. This more advanced choice is only used when we run PDLP without feasibility polishing.
Appendix B Derivation of Equation 3
This section shows how Equation 3 follows from specializing Algorithm 1 of Chambolle and Pock [27] to our saddle-point formulation
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}~\mathcal{L}(x,y)=c^{\top}x-y^{\top}Ax-p(y;-u_{c},-\ell_{c}).$
The standard saddle-point form of [27, Equation (2)] with $f(x)=0$ and finite-dimensional spaces is
$\min_{x\in{\mathbb{R}}^{n}}\max_{y\in{\mathbb{R}}^{m}}y^{\top}Kx+g(x)-h^{*}(y),$
where $g$ and $h$ are proper, lower semicontinuous, convex functions, and $h^{*}$ is the convex conjugate of $h$. This saddle-point problem reduces to our formulation by setting
$g(x):=c^{\top}x+I_{\mathcal{X}}(x),\qquad h^{*}(y):=I_{\mathcal{Y}}(y)+p(y;-u_{c},-\ell_{c}),\qquad K:=-A,$
where for any set $S$ and vector $s$ we define the indicator function
$I_{S}(s):=\begin{cases}\infty&\text{if }s\not\in S,\\ 0&\text{if }s\in S.\end{cases}$
The PDHG algorithm, i.e., Algorithm 1 of Chambolle and Pock [27] with the Euclidean norm, takes the form
$x^{k+1}$ $=\mathop{\rm argmin}_{x}g(x)+(y^{k})^{\top}Kx+\frac{1}{2\tau}\|x-x^{k}\|_{2}^{2},$ (11a)
$y^{k+1}$ $=\mathop{\rm argmin}_{y}h^{*}(y)-y^{\top}K(2x^{k+1}-x^{k})+\frac{1}{2\sigma}\|y-y^{k}\|_{2}^{2},$ (11b)
where $\tau,\sigma>0$ satisfy $\tau\sigma\|K\|_{2}^{2}<1$.
Primal step.
To see that Equation 11a reduces to Equation 3a, observe that
$x^{k+1}$ $=\mathop{\rm argmin}_{x}g(x)+(y^{k})^{\top}Kx+\frac{1}{2\tau}\|x-x^{k}\|_{2}^{2}=\mathop{\rm argmin}_{x}c^{\top}x+I_{\mathcal{X}}(x)+(y^{k})^{\top}(-A)x+\frac{1}{2\tau}\|x-x^{k}\|_{2}^{2}$
$=\mathop{\rm argmin}_{x\in\mathcal{X}}c^{\top}x-(y^{k})^{\top}Ax+\frac{1}{2\tau}\|x-x^{k}\|_{2}^{2}=\mathop{\rm argmin}_{x\in\mathcal{X}}(c-A^{\top}y^{k})^{\top}x+\frac{1}{2\tau}\|x-x^{k}\|_{2}^{2}$
$=\mathop{\bf proj}_{\mathcal{X}}(x^{k}-\tau(c-A^{\top}y^{k})).$
Dual step.
To see that Equation 11b reduces to Equation 3b, observe that
$y^{k+1}$ $=\mathop{\rm argmin}_{y}h^{*}(y)-y^{\top}K(2x^{k+1}-x^{k})+\frac{1}{2\sigma}\|y-y^{k}\|_{2}^{2}$
$=\mathop{\rm argmin}_{y}I_{\mathcal{Y}}(y)+p(y;-u_{c},-\ell_{c})-y^{\top}(-A)(2x^{k+1}-x^{k})+\frac{1}{2\sigma}\|y-y^{k}\|_{2}^{2}$
$=\mathop{\rm argmin}_{y\in\mathcal{Y}}p(y;-u_{c},-\ell_{c})+y^{\top}A(2x^{k+1}-x^{k})+\frac{1}{2\sigma}\|y-y^{k}\|_{2}^{2}$
$=\mathop{\rm argmin}_{y\in\mathcal{Y}}p(y;-u_{c},-\ell_{c})+\frac{1}{2\sigma}\|y-(y^{k}-\sigma A(2x^{k+1}-x^{k}))\|_{2}^{2}$
$=\mathbf{prox}_{\sigma p(\cdot;-u_{c},-\ell_{c})}(y^{k}-\sigma A(2x^{k+1}-x^{k})).$ (12)
The proximal operator of $\sigma p(\cdot;-u_{c},-\ell_{c})$ can be written as
$\mathbf{prox}_{\sigma p(\cdot;-u_{c},-\ell_{c})}(v)=v-\sigma\mathop{\bf proj}_{[-u_{c},-\ell_{c}]}(\sigma^{-1}v),$
due to Moreau’s decomposition, i.e., $v=\mathbf{prox}_{f}(v)+\mathbf{prox}_{f^{*}}(v)$ for $f(v)=\sigma p(v;-u_{c},-\ell_{c})$, and the fact that convex conjugate of $f$ is $f^{*}(v)=I_{[-u_{c},-\ell_{c}]}(\sigma^{-1}w)$. Substituting this expression into Equation 12 yields
$y^{k+1}$ $=y^{k}-\sigma A(2x^{k+1}-x^{k})-\sigma\mathop{\bf proj}_{[-u_{c},-\ell_{c}]}(\sigma^{-1}(y^{k}-\sigma A(2x^{k+1}-x^{k})))$
$=y^{k}-\sigma A(2x^{k+1}-x^{k})-\sigma\mathop{\bf proj}_{[-u_{c},-\ell_{c}]}(\sigma^{-1}y^{k}-A(2x^{k+1}-x^{k})),$
which matches Equation 3b.
Appendix C Theoretical basis for feasibility polishing
For simplicity, we focus on PDHG with a fixed step size less than $1/\|A\|_{2}$ and a fixed frequency restart scheme. The former choice is typical for analyzing PDHG [26, 27].
Also for simplicity, we assume the LP instance takes the form $\min c^{\top}x$ such that $Ax=b,x\geq 0$ and focus on primal feasibility polishing (i.e., $c=\mathbf{0}$). However, this analysis generalizes to other settings such as the LP format described in Section 2 and dual feasibility polishing. In this simple case, PDHG reduces to
@@ Appendix C Theoretical basis for feasibility polishing
Let $X_{F}:=\{x\geq 0:Ax=b\}$ be the set of feasible solutions. Let $\mathop{\bf dist{}}(X,x):=\inf_{z\in X}\|x-z\|_{2}$. The Hoffman constant of $A$ is the largest constant, $H(A)$, such that
@@
Thanks to Hoffman [55], we know that this constant is always positive.
@@
Let $Z^{\star}$ denote the set of minimax solutions to $\min_{x\geq 0}\max_{y}b^{\top}y-y^{\top}Ax$. From Corollary 2 of [10] we have
@@
Moreover, by Applegate et al. [10] equation (22), we have
@@
where ++ is the positive part operator. Combining (14), (17), (18), and (19) yields (16). ∎
Input: A starting point $x^{0,0}$, a step-size $\eta\in(0,\infty)$, a length for the restart period $\ell\in{\mathbb{N}}$;
Initialize outer loop counter $n\leftarrow 0$;
Initialize duals $y^{0,0}\leftarrow\mathbf{0}$;
for $n=0,1,2,\dots$ do
 initialize the inner loop. inner loop counter $t\leftarrow 0$ ;
 repeat
      Take a step of PDHG, i.e., calculate $(x^{n,t+1},y^{n,t+1})$ from $(x^{n,t},y^{n,t})$ via (13) ;
    $\bar{x}^{n,t+1}\leftarrow\frac{1}{t+1}\sum_{i=1}^{t+1}x^{n,i}$;
    $t\leftarrow t+1$;
 until $t\geq\ell$ ;
 restart the outer loop. $x^{n+1,0}\leftarrow\bar{x}^{n,t+1}$, $y^{n+1,0}\leftarrow\mathbf{0}$, $n\leftarrow n+1$;
end for
@@
$\|x^{k,0}-x^{0,0}\|_{2}$ $\leq\|x^{k,0}-x^{k-1,0}\|_{2}+\|x^{k-1,0}-x^{0,0}\|_{2}\leq(q+2)\mathop{\bf dist{}}(x^{k-1,0},X_{F})+\|x^{k-1,0}-x^{0,0}\|_{2}$ (21)
$\leq(q+2)2^{1-k}\mathop{\bf dist{}}(x^{0,0},X_{F})+\|x^{k-1,0}-x^{0,0}\|_{2}$
@@
$\|x^{n,0}-x^{0,0}\|_{2}\leq\left(\sum_{i=1}^{n}2^{1-i}\right)(q+2)\mathop{\bf dist{}}(x^{0,0},X_{F}).$
@@
In particular, (21) with $k=1$ shows the hypothesis for $n=1$ since $x^{k-1,0}=x^{0,0}$. If the hypothesis holds for n=k1n=k-1 then by (21) it holds for $n=k$. Thus we have proved the induction hypothesis. Using the formula for the sum of an infinite geometric series we get $\sum_{i=1}^{n}2^{1-i}\leq\sum_{i=1}^{\infty}2^{1-i}=2$, thus the induction hypothesis implies (20b). ∎
Equation 20a implies quick convergence of feasibility polishing to a feasible solution. Equation 20b implies feasibility polishing will converge to a feasible solution that is approximately optimal, assuming the starting point $x^{0,0}$ was approximately optimal. In particular, Equation 20b implies,
[…]
Moreover, taking $\eta=\alpha/\|A\|_{2}$ where $\alpha\in(0,1)$, as typical [26, 27], gives $q=4\frac{1+\alpha}{1-\alpha}$. For example, $\alpha=1/2$ gives $q=12$. Thus, from Equation 22, we conclude that if we start at a point $x^{0,0}$ with $c^{\top}x^{0,0}\approx c^{\top}x^{\star}$ and $\|Ax^{0,0}-b\|_{2}\approx 0$ then
[…]
Appendix D Infeasibility termination criteria
Our infeasibility termination criteria closely follow those of SCS [79] and Applegate et al. [11].
We declare the problem primal infeasible if we find a point $(y,r)\in\mathcal{Y}\times\mathcal{R}$ satisfying:
$p(-y;\ell_{c},u_{c})-p(-r;\ell_{v},u_{v})>0,\quad\frac{\|A^{\top}y+r\|_{\infty}}{-p(-y;\ell_{c},u_{c})-p(-r;\ell_{v},u_{v})}\leq 10^{-9}.$
Recall that
$p(y;\ell,u):=u^{\top}y^{+}-\ell^{\top}y^{-},$
and therefore
$p(-y;\ell_{c},u_{c})-p(-r;\ell_{v},u_{v})=\ell_{c}^{\top}y^{+}-u_{c}^{\top}y^{-}+\ell_{v}^{\top}r^{+}-u_{v}^{\top}r^{-}.$
We declare the problem dual infeasible if we find a point $(x,s)\in\bar{\mathcal{X}}\times\bar{\mathcal{S}}$ satisfying:
[…]
where $\bar{\mathcal{X}}\subseteq{\mathbb{R}}^{n}$ and $\bar{\mathcal{S}}\subseteq{\mathbb{R}}^{m}$ are Cartesian products whose $i$th components are given by
$\begin{aligned} &\bar{\mathcal{S}}_{i}:=\begin{cases}\{0\}&(\ell_{c})_{i},(u_{c})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{c})_{i}=-\infty,~(u_{c})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{-}&(\ell_{c})_{i}\in{\mathbb{R}},~(u_{c})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\quad\text{ and }\qquad\end{aligned}\begin{aligned} &\bar{\mathcal{X}}_{i}:=\begin{cases}\{0\}&(\ell_{v})_{i},(u_{v})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{v})_{i}\in{\mathbb{R}},(u_{v})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{v})_{i}=-\infty,(u_{v})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}&\text{otherwise}.\end{cases}\end{aligned}$
Remarks:
  • The threshold of $10^{-9}$ was selected to ensure that none of the large-scale test instances would incorrectly terminate with an infeasibility declaration.
  • All instances were rescaled prior to solving (Section 6), which implicitly affects the infeasibility termination criteria in a manner analogous to how rescaling affects the optimality termination criteria; see Equation 10.
Appendix E Detailed statistics for PDLP on the large-scale instances
Table 5: Detailed statistics for PDLP on the large-scale instances. KKT passes refers to the total number of times that $Ax$ and $A^{\top}y$ are computed. Residuals are measured in $\ell_{\infty}$-norm, i.e., absolute error (on the instances scaled as per Section 6). The gap is the difference between the primal and dual objective. Rows with gray background indicate unsuccessful termination (i.e., either hitting the six day time limit or in the case of qap-tho-150, a numerical error as per Table 3), in which case we report the statistics from the most recent recorded iteration. Memory is the maximum memory usage in GB according to SLURM. Note that finer grained information is available at http://www.oliverhinder.com/large-scale-lp-problems.
Time (hrs) KKT passes Objective Residuals Memory (GB)
Instance name Primal Dual Gap Primal Dual
With polishing
design-match 9.34 1.5E+4 2.87E+6 2.87E+6 5.5E-2 4.0E-11 0 135
heat-source-easy 59.97 1.9E+6 5.16E+7 5.16E+7 9.1E-4 9.9E-9 9.3E-9 18
[…]
mediterranean-shipping 32.62 1.1E+5 -1.07E+3 -1.07E+3 6.9E+0 7.8E-9 0 81
[…]
qap-wil-100 0.28 3.8E+3 2.21E+5 2.21E+5 1.1E+2 5.7E-9 0 22
supply-chain 18.90 6.1E+4 1.08E+8 1.08E+8 6.7E+3 5.6E-9 0 70
tsp-gaia-100m 21.06 7.8E+3 1.94E+9 1.94E+9 5.8E+4 7.6E-9 0 611
tsp-gaia-10m 2.99 7.9E+3 1.30E+8 1.30E+8 3.0E+3 6.6E-9 0 40
world-shipping 53.81 2.2E+5 -1.18E+5 -1.18E+5 3.7E+2 9.9E-9 0 91
Without polishing
design-match 67.95 1.1E+5 2.87E+6 2.87E+6 -8.7E-6 9.8E-9 0 136
[…]
qap-wil-100 44.35 6.4E+5 2.21E+5 2.21E+5 1.6E-4 6.6E-9 1.0E-8 20
[…]
Appendix F Additional numerical experiments
To supplement the experiments on large-scale linear programs in the body of the paper, we ran some additional experiments to verify the performance on the Mittelmann benchmark (Section F.1), a standard collection of slightly smaller instances, and on a set of infeasible instances (Section F.2). For consistency, PDLP was set up exactly the same as it was for the large-scale experiments, including keeping presolved disabled, rescaling the instances, and maintaining the same termination criteria. For these experiments, we used Gurobi version 12.0.3 instead of 11.0.2 (the version used for the large-scale experiments in Section 7) because at the time of running these experiments, the Pitt computing cluster we use had updated its Gurobi version.
F.1 Numerical results on Mittelmann’s benchmark instances
We compared PDLP with and without feasibility polishing on the Mittelmann test set. Our results are reported in Figure 3, which shows that feasibility polishing is never much slower than standard PDLP and can give orders of magnitude speed improvements, consistent with our large-scale numerical results in Section 7.
We also tested the impact of reducing the feasibility tolerance (both for the primal and dual). The default feasibility tolerance is $10^{-8}$ as specified in Equation 9. These results are reported in Figure 4, which shows that PDLP with feasibility polishing is insensitive to reductions in feasibility tolerance, while PDLP without feasibility polishing suffers significant performance degradation as the feasibility tolerance is reduced. Without polishing, one additional instance failed to solve at $10^{-10}$ feasibility tolerance, and a further eight additional instances failed at $10^{-12}$ feasibility tolerance. With polishing, one additional instance failed to solve at $10^{-10}$ feasibility tolerance, and a further four additional instances failed at $10^{-12}$ feasibility tolerance.
LABEL:table:mittelmann-results reports the results for PDLP (with and without polishing), and Gurobi (barrier, primal simplex, and dual simplex) on the Mittelmann benchmark instances. From this table it is clear that, with the exception of three instances, Gurobi is almost always faster. However, it is important to keep in mind that this CPU implementation of PDLP studied in this paper is not intended for tackling this scale of problem and a GPU implementation is likely more suitable for such a comparison.
F.1.1 Instances incorrectly declared infeasible
PDLP, both with and without polishing, fails to find an optimal solution in 14 of 49 instances. Two of these failures occur because the instance was declared infeasible by PDLP (both with and without feasibility polishing): pds-100 and a2864. We found that dropping the infeasibility termination tolerance to $10^{-10}$ prevents pds-100 from being declared infeasible, allowing the problem to be solved in 1075.1 seconds with feasibility polishing and in 1727.3 seconds without feasibility polishing. Decreasing the infeasibility termination tolerance for a2864 does not prevent it from being declared infeasible, and detailed investigation indicates that the issue is likely a small numerical error either in the computation of the infeasibility criteria or in the rescaling scheme (See https://github.com/google/or-tools/issues/4874). Manually modifying the infeasibility criteria from $p(-y;\ell_{c},u_{c})-p(-r;\ell_{v},u_{v})>0$ to $p(-y;\ell_{c},u_{c})-p(-r;\ell_{v},u_{v})>10^{-14}$ fixes the issue, allowing the problem to be solved in 116.4 seconds with feasibility polishing and in 351.9 seconds without feasibility polishing.
Figure 3: Comparison of PDLP with and without feasibility polishing for the Mittelmann test set. This test set consists of 49 problems, 14 of which were unsolved by both with and without polishing. Note that all problems solved without polishing were also solved with the polishing option enabled.
Figure 4: Comparison of PDLP with and without feasibility polishing for the Mittelmann test set with different feasibility tolerances (both for the primal and dual). The standard feasibility tolerance in this paper is $10^{-8}$.
Table 6: Statistics for Mittelmann’s linear programming collection. Instances collected are based on the benchmark pages from 2025-06-18, which can be retrieved from archive.org: https://web.archive.org/web/20250624111618/https://plato.asu.edu/ftp/lpfeas.html.
problem # constraints # variables # nonzeros
L1_sixm250obs 986069 428032 4280320
Linf_520c 93326 69004 566193
a2864 22117 200787 20078717
bdry2 376500 250998 1500003
cont1 160793 40398 399991
cont11 160793 80396 439989
datt256 11077 262144 1503732
dlr1 1735470 9121907 18365107
ex10 69609 17680 1179680
fhnw-bin1 772872 1141653 8611326
fome13 48569 97840 334984
graph40-40 360900 102600 1260900
irish-e 104260 61728 538809
neos 479120 36786 1084461
neos3 512209 6624 1542816
neos-3025225 91572 69846 9357951
neos5052403 38269 32868 4898304
neos-5251015 486531 136971 1955388
ns1687037 50622 43749 1406739
ns1688926 32768 16587 1712128
nug08-3rd 19728 20448 139008
pds-100 156244 505360 1390539
psched3-3 266228 79555 1062480
qap15 6331 22275 110700
rail02 95791 270869 756228
rail4284 4284 1092610 12372358
rmine15 358395 42438 879732
s82 87878 1690631 7022608
s100 14734 364417 2127672
s250r10 10963 273142 1572104
savsched1 295990 328575 1846351
scpm1 5000 500000 6250000
shs1023 133944 444625 1044725
square41 40161 62234 13628623
stat96v2 29089 957432 2852184
stormG2_1000 528186 1259121 4228817
stp3d 159488 204880 662128
support10 165685 14770 551152
tpl-tub-ws16 1154615 747691 4720567
woodlands09 194599 382147 2646003
Dual2_5000 30000600 33050602 93001800
Primal2_1000 1299380 2559380 5498140
thk_48 6366377 8609262 27802878
thk_63 5694387 7701112 21592414
L1_sixm1000obs 3082940 1426256 14262560
L2CTA3D 210000 10000000 30000000
degme 185501 659415 8127528
dlr2 7132926 38868107 78091589
set-cover 10000 1102008 20442268
[…]
Problem PDLP Gurobi (minutes)
KKT Passes Time (minutes) Barrier Simplex
no polish polish no polish polish Primal Dual
Dual2_5000 79,497 79,497 193.36 194.59 $\dagger$ 59.19
L1_sixm1000obs 69,577 7,049 35.19 3.55 1.46 2.18 0.53
L1_sixm250obs 53,324 6,733 12.53 1.70 0.11 0.11 0.09
L2CTA3D 19,872 987 13.16 1.10 22.03
Linf_520c 0.09 9.20 1.00
Primal2_1000 922,873 293.87 1181.92 1236.06
a2864 $\times$ $\times$ $\times$ $\times$ 0.03 5.06 6.20
bdry2 0.27 71.18 96.51
cont1 3,331,570 618.30 0.03 5.13 5.89
cont11 3,693,319 714.41 0.04 38.09 25.66
datt256 1,347 1,499 0.27 0.33 0.03 225.96 5.28
degme 0.73
dlr1 0.91 210.87 39.98
dlr2 9.14
ex10 710 788 0.17 0.20 0.06 0.08
fhnw-bin1 357,434 357,594 108.48 109.10 2.54 0.45 3.86
fome13 440,673 84.78 0.02 0.12 0.13
graph40-40 1,090 1,286 0.21 0.29 0.02 0.39 0.12
irish-e 0.05 0.74 0.39
neos 4,054,159 794.93 0.15 1.87 0.11
neos-3025225 11,141 11,141 3.02 3.80 0.21 7.83 2.55
neos-5251015 2,805 2,805 0.53 0.54 0.06 47.87 0.05
neos3 854 822 0.18 0.17 0.03 0.43 0.05
neos5052403 40,302 1,700 9.51 0.46 0.08 2.83 0.85
ns1687037 0.08 10.56 4.58
ns1688926 0.03 1.16 0.02
nug08-3rd 1,537 1,689 0.34 0.53 0.01 0.82 0.62
pds-100 $\times$ $\times$ $\times$ $\times$ 0.25 1.97 0.17
psched3-3 0.13 1.51 1.74
qap15 50,434 1,926 9.81 0.55 0.01 0.21 0.30
rail02 0.69 52.74 39.41
rail4284 219,460 960 65.48 0.39 0.32 30.54 0.80
rmine15 70,824 7,834 14.26 1.70 0.25 10.29 8.79
s100 117,222 105,378 20.99 19.60 0.12 1.33 1.71
s250r10 2,812,856 2,864,646 505.42 502.39 0.11 1.41 0.32
s82 2,082,799 1,808,861 780.22 642.19 1.26 165.40 20.05
savsched1 1,955 966 0.40 0.26 0.13 0.46 1.74
scpm1 1,815 1,639 0.53 0.51 0.25 13.33 4.54
set-cover 688,191 3,385 310.96 1.80 1.69 755.10 1041.93
shs1023 0.29 17.77 1.12
square41 242,433 232,357 60.14 59.14 0.03 0.06 0.05
stat96v2 0.38 108.24 30.46
stormG2_1000 48,713 30,365 12.90 7.98 0.58 13.71 0.38
stp3d 92,890 15,187 16.48 2.89 0.07 10.50 1.39
support10 21,540 6,736 3.96 1.25 0.06 0.69 0.28
thk_48 140,512 14,658 122.21 13.20 58.65
thk_63 32,774 29,894 28.45 24.06 2.18 12.94
tpl-tub-ws16 713,992 208,841 236.37 69.26 0.54 6.00 13.90
woodlands09 17,870 17,870 3.89 3.72 0.08 4.77 8.32
F.2 Numerical results for primal infeasibility detection
We compared PDLP with and without feasibility polishing on the NETLIB infeasibility test set (See https://www.netlib.org/lp/infeas/readme). This test set mostly comprises smaller instances (See LABEL:table:infeasibility-problem-stats), but it can be a good test of robustness.
PDLP performance on this set is reported in LABEL:table:infeasibility-performance. Overall, PDLP’s performance on these problems is less than desired, with it only solving 15/29 problems. Thus, improving the ability of first-order methods to reliably detect infeasibility is an important subject for future research. We also find that on these instances, feasibility polishing has no notable impact on solver performance.
We also tested Gurobi on these problems but do not report the results because it was so dominant. In particular, every instance was declared primal infeasible by at least one Gurobi method in less than 0.1 seconds, except ‘pang’, which was declared either infeasible or unbounded by both dual and primal simplex in 0.01 seconds.
Table 8: Problem statistics for primal infeasible linear programs.
problem # constraints # variables # nonzeros
bgdbg1 348 407 1440
bgetam 400 688 2409
bgindy 2671 10116 65502
bgprtr 20 34 64
box1 231 261 651
ceria3d 3576 824 17602
chemcom 288 720 1566
cplex1 3005 3221 8944
cplex2 224 221 1058
ex72a 197 215 467
ex73a 193 211 457
forest6 66 95 210
galenet 8 8 16
gosh 3792 10733 97231
gran 2658 2520 20106
greenbea 2393 5405 30883
itest2 9 4 17
itest6 11 8 20
klein1 54 54 696
klein2 477 54 4585
klein3 994 88 12107
mondou2 312 604 1208
pang 361 460 2652
pilot4i 410 1000 5141
qual 323 464 1646
reactor 318 637 2420
refinery 323 464 1626
vol1 323 464 1646
woodinfe 35 89 140
Table 9: Computational results for the collection of primal infeasible linear programs given in LABEL:table:infeasibility-problem-stats. The time limit for these problems was one hour. One KKT pass refers to the calculation AxAx and $A^{\top}y$.
Problem Time (seconds) KKT Passes
no polishing polishing no polishing polishing
bgdbg1 104.0 87.6 7379 7379
bgetam 239.9 283.7 22160 22160
bgindy 1.9 2.0 68 68
bgprtr 14.3 15.6 1857 1857
box1 139.0 142.0 11973 11973
ceria3d
chemcom 34.7 36.2 3081 3081
cplex1
cplex2
ex72a 360.6 380.7 31990 31990
ex73a 438.1 554.3 41611 41623
forest6 64.3 66.6 5570 5570
galenet 0.1 0.1 67 67
gosh
gran
greenbea 231.4 214.5 18131 18131
itest2 0.1 0.1 206 206
itest6 0.1 0.1 66 66
klein1
klein2
klein3
mondou2 113.7 114.2 7693 7693
pang
pilot4i
qual
reactor
refinery
vol1
woodinfe 41.4 37.8 3969 3969